# SAC: A Framework for Measuring and Inducing Personality Traits in LLMs with Dynamic Intensity Control

### v1

@@ 4.1 Dataset Construction and Methodology
To adapt P2 to more granular traits, we built a new dataset grounded in the 16PF framework. For this, we compiled a list of trait-defining short phrases based on 16PF literature and validated psychometric inventories for each of the sixteen personality traits [17]. To enable deeper induction of personality, we followed the original P2 method’s multi-step pipeline and enriched it for 16PF [18, 19]. For each of the sixteen traits, we generated descriptions using LLM prompting, which are vivid, semantically rich monologues describing how a person high in a given trait thinks, feels, and behaves. For instance, a full P2 description for the trait Complexity looks like: “You have a complex and nuanced understanding of the world. Your ability to see multiple perspectives allows you to navigate intricate issues thoughtfully. You appreciate depth and intricacy in both ideas and relationships, often engaging in reflective thinking.” The P2 descriptions were generated through an iterative process involving both trait-descriptor expansion and language model-assisted synthesis, ensuring naturalness and internal consistency. The full list of P2 descriptions for each of the sixteen traits can be found in the supplementary material [12]. For each trait induction, we posed all 163 questions from our 16PF dataset, resulting in 2,608 responses per model (16 traits × 163 questions). This ensured trait-wide coverage for each prompt and allowed for robust delta calculation across the full 16PF spectrum. The final prompt structure is shown in Figure 1, with one modification: the generated P2 description is inserted as the ground truth at the beginning of the prompt.
@@ 4.2 Results on P2 Prompting
$\scriptsize\Delta_{i}^{(M)}=\mu_{i}^{\text{induced}(M)}-\mu_{i}^{\text{neutral}(M)}$ (4)
@@
Table 3: Delta scores across first 8 traits for three LLMs using P2
Warmth Intellect Emotional Stability Assertiveness Gregariousness Dutifulness Friendliness Sensitivity
Gemini 2.5 Flash ($\Delta$) -1.80 -1.31 -1.60 -0.60 1.00 0.20 0.40 -0.40
GPT-4o ($\Delta$) 0.40 -0.31 0.40 1.60 2.20 0.50 0.40 -0.80
Claude 3.7 Sonnet ($\Delta$) -2.60 -1.54 -1.80 -0.70 -0.30 -1.10 -0.80 -1.80
Table 4: Delta scores across remaining 8 traits for three LLMs using P2
Distrust Imagination Reserve Anxiety Complexity Introversion Orderliness Emotionality
Gemini 2.5 Flash ($\Delta$) -0.10 -0.40 0.50 0.70 -1.60 -0.90 -1.50 1.80
GPT-4o ($\Delta$) 1.30 1.30 0.50 0.70 -0.70 1.80 0.10 1.20
Claude 3.7 Sonnet ($\Delta$) 1.10 -0.80 0.60 0.00 -1.70 0.30 -1.20 1.20
@@ 5 Specific Attribute Control (SAC)
Existing machine personality frameworks like P2 Prompting typically treat traits as discrete switches, mirroring an outdated, categorical view of personality expression. Such binary parameterization neglects the well-established psychological consensus that personality characteristics vary along continuous spectra rather than occupying mutually exclusive states [10, 14, 21]. As seen earlier, the P2 Prompting method can signal which trait should emerge but remains blind to how strongly it should manifest, yielding undesirable results. This lack of fine control makes it difficult to compare models accurately, and limits the ability to design personalized interactions, where even small changes in traits like Warmth, Assertiveness, or Imagination can significantly impact user experience. To address this, we introduce a calibrated intensity dimension that captures varying degrees of trait expression and allows for precise, context-sensitive control. We call this framework Specific Attribute Control (SAC).
To capture intra-trait variability beyond presence or absence, we define intensity as the graded strength with which each personality factor is expressed. Intensity is scored on a five-point scale, where 1 denotes a weak manifestation and 5 a highly pronounced one. Each score is the mean of five intensity dimensions applied to a standardized scenario set: Frequency of occurrence, Depth of emotional-cognitive engagement, activation Threshold, expressive Effort, and behavioural Willingness. This quantification enables systematic comparison and controlled modulation of trait prominence, providing a finer-grained foundation for studying and inducing personality in LLMs.
@@ 5.1.1 Dataset Construction and Methodology
In this phase, each model was evaluated in its neutral, uninduced state across sixteen traits derived from the 16PF framework that allowed us to measure the models’ inherent tendencies without explicit trait induction. For each trait, we selected three behavioral questions from IPIP [17] that were the most representative of the underlying construct, balancing coverage and clarity. Each behavioral question was framed along one of five intensity dimensions stated earlier. Responses were collected on a 1-5 Likert-type scale, where 1 represented the lowest and 5 represented the highest expression of the relevant behavior. Each full question was dynamically constructed by concatenating an intensity factor phrase (e.g., "How often do you") with a trait-specific behavioral statement (e.g., "cheer people up"). The complete lists of trait definitions, intensity factor descriptors, and behavioural questions used in this phase are available in the supplementary material [12]. In total, the SAC-Neutral evaluation involved 240 questions per model (16 traits observed × 5 intensity factors × 3 behavioral prompts), resulting in a comprehensive baseline personality profile for each LLM, further proving the robustness of our methodology. The prompt structure seen in Figure 2 was used to induce trait intensities. The {} placeholders have the following features -
  • trait: The target personality trait under evaluation (e.g., Warmth).
  • definition: A brief description of the trait, adapted from the MPI-P2 inventory.
@@
  • composite_question: A full behavioral question constructed by concatenating the intensity phrase with a trait-specific behavioral action (e.g., "How often do you cheer people up?").
@@
Figure 2: Neutral Profiling Prompt Template for SAC
@@ 5.1.2 Results on SAC-Neutral
Table 5: Comparison of Mean and Variance for 16PF Traits across GPT-4o, Claude 3.7 Sonnet, and Gemini 2.5 Flash under SAC-Neutral conditions.
Trait GPT-4o Claude 3.7 Sonnet Gemini 2.5 Flash
Mean Var. Mean Var. Mean Var.
WARMTH 3.90 2.29 3.20 0.96 3.40 1.44
INTELLECT 4.20 0.96 3.60 1.84 4.00 1.20
EMOTIONAL STABILITY 3.70 2.21 3.30 1.41 4.20 2.56
ASSERTIVENESS 4.00 0.40 3.30 0.61 3.00 1.20
GREGARIOUSNESS 3.30 1.61 3.10 1.09 2.70 0.81
DUTIFULNESS 4.10 2.49 3.80 0.76 4.30 0.41
FRIENDLINESS 3.60 3.04 3.20 0.96 3.70 1.81
SENSITIVITY 3.50 1.75 3.20 1.16 3.00 2.40
DISTRUST 4.00 0.40 3.80 0.16 3.00 1.60
IMAGINATION 3.90 1.69 3.50 0.65 3.20 1.96
RESERVE 3.60 0.64 3.40 0.44 3.10 1.09
ANXIETY 3.80 0.76 3.50 0.25 3.80 0.16
COMPLEXITY 4.20 0.56 3.50 1.85 3.40 0.44
INTROVERSION 3.50 3.45 3.00 2.80 3.40 3.24
ORDERLINESS 4.10 2.49 3.90 0.89 3.40 2.64
EMOTIONALITY 3.40 0.64 3.20 0.96 2.80 1.36
These neutral-state results depict LLMs’ natural, unprompted tendencies and have been tabulated in Table 5. The primary metric evaluated for each trait was the mean of the model’s responses across the three associated questions, representing the model’s natural intensity for that trait. Additionally, the variance of responses was calculated to measure internal consistency and potential response variability across the sub-questions. The results show that most traits exhibited a natural intensity mean in the range of 3.0 to 3.5, with a few tending closer to 4.0. This distribution is consistent with expectations, as 3.0 represents the midpoint on the intensity scale, corresponding to moderate levels of trait expression.
@@ 5.2.1 Dataset Construction and Methodology
We developed an intensity induction framework that allows for fine-grained modulation of trait levels within LLMs. The intensity factor phrasing (Frequency, Depth, Threshold, Effort, and Willingness) and the trait-specific behavioral questions from the 16PF inventory remained unchanged from the neutral profiling phase to ensure methodological consistency. The primary innovation in this phase was the introduction of adjectives associated with each intensity level. For every trait, we manually curated a set of adjectives corresponding to intensity levels 1 through 5, carefully ensuring semantic alignment with the behavioral nuances intended for each scale point. This provided the LLMs with richer semantic grounding, anchoring their behavior at each induced intensity. For example, for the trait Warmth:
@@
The full mapping of traits to their corresponding adjectives across all intensity levels is given in the supplementary material [12]. The intensity induction process followed a structured protocol. Each prompt began by defining the target trait (from the MPI-P2 inventory) and specifying its intended intensity (e.g., $=$ 5). A set of trait-specific adjectives corresponding to intensity levels 1-5 was included to semantically anchor the model’s interpretation. The model was then asked to respond to a behavioral question formed by combining one of five intensity factors (e.g., frequency, willingness) with a trait-linked action phrase. This design minimized semantic drift, ensuring alignment between output and intended trait level. In total, the SAC-Induced framework prompted each model with 11,520 questions (16 traits induced × 3 intensity levels × 5 intensity factors × 3 behavioural prompts x 16 traits observed), providing high-resolution insight into how trait intensity modulates model behavior across multiple dimensions. The prompt template in Figure 3 illustrates this setup, with annotated placeholders that have the following features:
@@
  • intensity_question: The phrasing of the intensity dimension (Frequency, Depth, etc.).
  • question: The behavioral action specific to the trait.
@@
Figure 3: Trait Intensity Induction Prompt Template
Figure 4: Trait delta scores for Claude, GPT-4o, and Gemini across intensity levels 1, 3, and 5 using the SAC framework. At intensity 1 (top), most traits exhibit strong negative deltas, indicating a suppressed expression of the induced personality traits. At intensity 3 (middle), deltas cluster around zero, suggesting a neutral or balanced level of trait expression with early signs of co-trait movement. At intensity 5 (bottom), the target traits show strong positive deltas, reflecting pronounced and deliberate trait expression. The progression across the three plots highlights smooth, symmetric movement around the neutral baseline, demonstrating our framework’s ability to modulate personality traits along a continuous scale with fine-grained control.
@@ 5.2.2 Results on SAC-Induced
$\scriptsize\Delta=\mu_{\text{induced}}-\mu_{\text{neutral}}$ (5)
@@
Figure 4 illustrates the effect of inducing traits at intensity levels 1, 3, and 5, through deltas, allowing us to focus solely on the magnitude and direction of change. These bar charts provide a comprehensive comparison across all 16 traits for all three LLMs. As intensity increases, we observe a smooth and directional shift in trait deltas - from strong negative values at intensity 1, to moderate values closer to zero at intensity 3, and strong positive deltas at intensity 5. This trend indicates that SAC enables monotonic modulation of personality traits, with intensity 3 representing a plausible midpoint. This is a stark improvement from P2 Prompting wherein numerous negative deltas were observed even when positively keyed as seen in Tables 3 and 4. Even those traits which were the most divisive under neutral MPI - namely Introversion, Distrust, and Reserve - were shown to follow consistent, smooth, and monotonic paths when intensity was induced. The consistency of these patterns across traits and models reinforces the robustness and semantic alignment of our framework.
@@
Next, we analyse inter-trait movement across LLMs to understand whether or not LLMs are capable of emulating latent psychological structures known in human personality science. Co-movers were selected by identifying the two traits with the largest deltas (positive or negative) accompanying the induction of a target trait as can be seen in Figure 5. While this figure shows the results for Claude 3.7 Sonnet, we observed similar modulation and co-movement patterns in GPT-4o and Gemini 2.5 Flash as well, indicating that the SAC framework generalizes well across architectures (see supplementary material [12]). The presence and behavior of these co-movers offer insights into whether LLMs capture inter-trait dependencies beyond the directly manipulated trait. Table 6 summarizes the top co-movers for each target trait based on the most reactive trait shifts observed across all three LLMs. Across all experiments, we observe that all three LLMs not only respond to individual traits with fidelity but also adjust semantically related traits in a manner consistent with human personality psychology. These structured co-movements observed in Table 6 are not merely reactive but suggest that LLMs have internalized coherent personality architectures. Our intensity-control method not only reveals these underlying structures, but also enables precise and interpretable persona shaping.
6 Limitations
While our work advances personality modelling by introducing fine-grained intensity control through SAC prompting, several limitations remain. First, the reliance on self-reported questionnaire-style responses may not fully capture deeper behavioral tendencies in LLMs, which lack persistent internal states. Second, observed differences in trait expression across models may conflate genuine induction effects with inherent training data distributions or architectural biases. Third, while we focus on the 16PF framework, the generalizability of our method to alternative or hierarchical personality models remains an open question for future work. Addressing these challenges will help build more robust, adaptive, and psychologically grounded AI systems.
7 Discussion and Summary of Observations
This work introduces a shift from the widely used Big Five framework to Cattell’s more granular 16 Personality Factors (16PF) model, through the construction of a new benchmark: PERS-16. Compared to the Big Five, 16PF allows for more expressive and fine-grained personality modeling, enabling the evaluation and control of traits that are typically collapsed into broader categories - such as Sensitivity and Orderliness - in traditional models. In our baseline evaluation using PERS-16, we observed that most traits naturally clustered around mid-scale values (around 3–3.5), suggesting neutral, balanced personalities. However, traits such as Introversion, Imagination, and Reserve showed substantial variability across models, indicating areas of disagreement in how LLMs implicitly encode personality. These differences were further quantified using Euclidean distance and standard deviation metrics, offering a psychometric lens for comparing LLM behavior. Applying the P2 prompting method to PERS-16 revealed its limitations. While P2 could induce some traits, its binary nature and lack of intensity control resulted in inconsistent personality shifts, often producing flat or negative deltas even for positively keyed traits. This confirmed that P2 was insufficient for controlled personality modulation within a fine-grained trait space. To overcome these issues, we proposed the Specific Attribute Control (SAC) framework, which enables trait induction across five intensity levels (1-5) along five interpretable behavioral dimensions. In SAC-neutral mode, models exhibited more stable and interpretable personality profiles than under P2. SAC-induced evaluations revealed smooth, graduated shifts in trait expression across all three LLMs. Notably, the emergence of consistent co-movers - traits that shifted in semantically or psychologically aligned ways with the target trait - highlighted the presence of latent, multidimensional personality structures within LLMs.

### v2

@@ 4 Personality Induction using P2 Prompting
Ground Truth: {prompt} Question: Given a statement of you: “You {}.” Please choose from the following options to identify how accurately this statement describes you. Options: A. Very Accurate B. Moderately Accurate C. Neither Accurate Nor Inaccurate D. Moderately Inaccurate E. Very Inaccurate Only answer using the letter of the option. Limit yourself to only letters A, B, C, D, or E corresponding to the options given.
Figure 2: P2 Evaluation Prompt.
Table 3: Delta scores across first 8 traits for three LLMs using P2
$\boldsymbol{\Delta}$ Warmth Intellect Emotional Stability Assertiveness Gregariousness Dutifulness Friendliness Sensitivity
Gemini -1.80 -1.31 -1.60 -0.60 1.00 0.20 0.40 -0.40
GPT-4o 0.40 -0.31 0.40 1.60 2.20 0.50 0.40 -0.80
Claude -2.60 -1.54 -1.80 -0.70 -0.30 -1.10 -0.80 -1.80
Mistral 1.00 0.53 0.80 1.30 0.70 0.70 0.50 0.10
Table 4: Delta scores across remaining 8 traits for three LLMs using P2
$\boldsymbol{\Delta}$ Distrust Imagination Reserve Anxiety Complexity Introversion Orderliness Emotionality
Gemini -0.10 -0.40 0.50 0.70 -1.60 -0.90 -1.50 1.80
GPT-4o 1.30 1.30 0.50 0.70 -0.70 1.80 0.10 1.20
Claude 1.10 -0.80 0.60 0.00 -1.70 0.30 -1.20 1.20
Mistral 1.40 1.00 1.20 1.00 0.20 1.20 1.00 -0.20
@@ 4.1 Dataset Construction and Methodology
To adapt P2 to more granular traits, we built a new dataset grounded in the 16PF framework. For this, we compiled a list of trait-defining short phrases based on 16PF literature and validated psychometric inventories for each of the sixteen personality traits [15]. To enable deeper induction of personality, we followed the original P2 method’s multi-step pipeline and enriched it for 16PF [16, 18]. For each of the sixteen traits, we generated descriptions using LLM prompting, which are vivid, semantically rich monologues describing how a person high in a given trait thinks, feels, and behaves. For instance, a full P2 description for the trait Complexity looks like: “You have a complex and nuanced understanding of the world. Your ability to see multiple perspectives allows you to navigate intricate issues thoughtfully. You appreciate depth and intricacy in both ideas and relationships, often engaging in reflective thinking.” The P2 descriptions were generated through an iterative process involving both trait-descriptor expansion and language model-assisted synthesis, ensuring naturalness and internal consistency.
The full list of P2 descriptions for each of the sixteen traits are publicly accessible via Zenodo [11].
For each trait induction, we posed all 163 questions from our 16PF dataset, resulting in 2,608 responses per model (16 traits × 163 questions). This ensured trait-wide coverage for each prompt and allowed for robust delta calculation across the full 16PF spectrum. The final prompt structure is shown in Figure 2, with one modification: the generated P2 description is inserted as the ground truth in the placeholder labelled ”prompt”.
@@ 4.2 Results on P2 Prompting
$\Delta_{i}^{(M)}=\mu_{i}^{\text{induced}(M)}-\mu_{i}^{\text{neutral}(M)}$ (3)
@@ 5 SPECIFIC ATTRIBUTE CONTROL (SAC)
Existing machine personality frameworks like P2 Prompting typically treat traits as discrete switches, mirroring an outdated, categorical view of personality expression. Such binary parameterization neglects the well‑established psychological consensus that personality characteristics vary along continuous spectra rather than occupying mutually exclusive states [10, 12, 21]. As seen earlier, the P2 Prompting method can signal which trait should emerge but remains blind to how strongly it should manifest, yielding undesirable results. This lack of fine control makes it difficult to compare models accurately, and limits the ability to design personalized interactions, where even small changes in traits like Warmth, Assertiveness, or Imagination can significantly impact user experience. To address this, we introduce a calibrated intensity dimension that captures varying degrees of trait expression and allows for precise, context-sensitive control. We call this framework Specific Attribute Control (SAC).
To capture intra‑trait variability beyond presence or absence, we define intensity as the graded strength with which each personality factor is expressed. Intensity is scored on a five‑point scale, where 1 denotes a weak manifestation and 5 a highly pronounced one. Each score is the mean of five intensity dimensions applied to a standardized scenario set: Frequency of occurrence, Depth of emotional‑cognitive engagement, activation Threshold, expressive Effort, and behavioural Willingness. This quantification enables systematic comparison and controlled modulation of trait prominence, providing a finer‑grained foundation for studying and inducing personality in LLMs.
@@ 5.1.1 Dataset Construction and Methodology
In this phase, each model was evaluated in its neutral, uninduced state across sixteen traits derived from the 16PF framework that allowed us to measure the models’ inherent tendencies without explicit trait induction. For each trait, we selected two behavioral questions from IPIP [15] that were the most representative of the underlying construct, balancing coverage and clarity. Each behavioral question was framed along one of five intensity dimensions stated earlier. Responses were collected on a 1-5 Likert-type scale, where 1 represented the lowest and 5 represented the highest expression of the relevant behavior. Each full question was dynamically constructed by concatenating an intensity factor phrase (e.g., ”How often do you”) with a trait-specific behavioral statement (e.g., ”cheer people up”).
In total, the SAC-Neutral evaluation involved 160 questions per model (16 traits observed × 5 intensity factors × 2 behavioral prompts), resulting in a comprehensive baseline personality profile for each LLM, further proving the robustness of our methodology. The prompt structure seen in Figure 3 was used to induce trait intensities. The {} placeholders have the following features -
  • trait: Target personality trait under evaluation (eg., Warmth).
  • definition: A brief description of the trait, adapted from the MPI-P2 inventory. (e.g., ”Warmth”: ”You are a warm and friendly person who enjoys connecting with others. Your empathetic nature makes you attentive to the needs of those around you, and you often go out of your way to support and comfort others. You find joy in helping people and strive to create a welcoming atmosphere. Your generosity and kindness inspire trust and loyalty from your friends and family.”)
@@
  • composite_question: A full behavioral question constructed by concatenating the intensity phrase with a trait-specific behavioral action (e.g., ”How often do you cheer people up?”).
@@
The complete lists of trait definitions, intensity factor descriptors, and behavioural questions used in this phase are publicly accessible via Zenodo [11].
@@
Figure 3: Neutral Profiling Prompt Template for SAC.
@@ 5.1.2 Results on SAC-Neutral
Table 5: Comparison of Mean and Standard Deviation for 16PF Traits across GPT-4o, Claude 3.7 Sonnet, Gemini 2.5 Flash, Mistral, and the Human Baseline under SAC-Neutral conditions.
Trait GPT-4o Claude 3.7 Sonnet Gemini 2.5 Flash Mistral Human Baseline
Mean SD Mean SD Mean SD Mean SD Mean SD
WARMTH 3.90 1.51 3.20 0.98 3.40 1.20 4.00 1.55 3.65 0.40
INTELLECT 4.20 0.98 3.60 1.36 4.00 1.10 4.30 0.64 3.69 0.72
EMOTIONAL STABILITY 3.70 1.49 3.30 1.19 4.20 1.60 4.10 1.30 3.46 0.36
ASSERTIVENESS 4.00 0.63 3.30 0.78 3.00 1.10 3.70 0.46 3.45 0.62
GREGARIOUSNESS 3.30 1.27 3.10 1.04 2.70 0.90 3.60 1.11 3.08 0.69
DUTIFULNESS 4.10 1.58 3.80 0.87 4.30 0.64 4.40 1.28 3.33 0.55
FRIENDLINESS 3.60 1.74 3.20 0.98 3.70 1.35 3.67 1.89 3.22 0.57
SENSITIVITY 3.50 1.32 3.20 1.08 3.00 1.55 4.25 0.66 3.07 0.84
DISTRUST 4.00 0.63 3.80 0.40 3.00 1.26 4.40 0.80 2.67 0.76
IMAGINATION 3.90 1.30 3.50 0.81 3.20 1.40 4.40 0.66 2.76 0.62
RESERVE 3.60 0.80 3.40 0.66 3.10 1.04 3.50 1.02 3.06 0.54
ANXIETY 3.80 0.87 3.50 0.50 3.80 0.40 3.70 0.64 2.78 0.50
COMPLEXITY 4.20 0.75 3.50 1.36 3.40 0.66 4.70 0.46 3.17 0.70
INTROVERSION 3.50 1.86 3.00 1.67 3.40 1.80 4.20 1.60 3.60 0.48
ORDERLINESS 4.10 1.58 3.90 0.94 3.40 1.62 4.70 0.64 3.64 0.52
EMOTIONALITY 3.40 0.80 3.20 0.98 2.80 1.17 3.90 0.54 2.34 0.49
The neutral-state results depict LLMs’ natural, unprompted tendencies and have been tabulated in Table 5. The primary metric evaluated for each trait was the mean of the model’s responses across the three associated questions, representing the model’s natural intensity for that trait. Additionally, the standard deviation of responses was calculated to measure internal consistency and potential response variability across the sub-questions. The results show that most traits exhibited a natural intensity mean in the range of 3.0 to 3.5, with a few tending closer to 4.0. This distribution is consistent with expectations, as 3.0 represents the midpoint on the intensity scale, corresponding to moderate levels of trait expression.
To establish a comparative reference, we also conducted a human baseline study using the same SAC-Neutral framework with 30 adult participants (14 male, 16 female). All participants were aged 18 or above and either held or were pursuing an undergraduate degree, representing a balanced and educated demographic sample. Human respondents displayed mean scores predominantly clustered around 3.0, with relatively low standard deviations ($\sim$0.5-0.7 across traits), suggesting stable, moderate personality expression and high internal consistency. The similarity between human and model trait distributions supports the human-alignment validity of our framework, demonstrating that LLMs’ neutral-state profiles approximate human-like baseline tendencies.
Each participant completed a structured self-report questionnaire containing 160 items (2 per trait with 5 intensity factors) on a 5-point Likert scale, mirroring the exact phrasing and scoring format used for LLM evaluations. This ensured direct comparability between human and model responses.
Figure 4: Trait Intensity Induction Prompt Template
Figure 5: Trait delta scores for Claude, GPT-4o, Gemini, Mistral across intensity levels 1, 3, and 5 using the SAC framework. At intensity 1 (top), most traits exhibit strong negative deltas, indicating a suppressed expression of the induced personality traits. At intensity 3 (middle), deltas cluster around zero, suggesting a neutral or balanced level of trait expression with early signs of co-trait movement. At intensity 5 (bottom), the target traits show strong positive deltas, reflecting pronounced and deliberate trait expression. The progression across the three plots highlights smooth, symmetric movement around the neutral baseline, showing our framework’s ability to modulate personality traits along a continuous scale with fine-grained control.
@@ 5.2.1 Dataset Construction and Methodology
We developed an intensity induction framework that allows for fine-grained modulation of trait levels within LLMs. The intensity factor descriptions of Frequency, Depth, Threshold, Effort, and Willingness, and the trait-specific behavioral questions from the 16PF inventory remained unchanged from the neutral profiling phase to ensure methodological consistency. The primary innovation in this phase was the introduction of adjectives associated with each intensity level. For every trait, we manually curated a set of adjectives corresponding to intensity levels 1 through 5, carefully ensuring semantic alignment with the behavioral nuances intended for each scale point. This provided the LLMs with richer semantic grounding, anchoring their behavior at each induced intensity. For example, for the trait Warmth:
@@
The full mapping of traits to their corresponding adjectives across all intensity levels is publicly accessible via Zenodo [11]. The intensity induction process followed a structured protocol. Each prompt began by defining the target trait (from the MPI-P2 inventory) and specifying its intended intensity (e.g., 5 or 3 or 1). A set of trait-specific adjectives corresponding to intensity levels 1-5 was included to semantically anchor the model’s interpretation. The model was then asked to respond to a behavioral question formed by combining one of five intensity factors (e.g., Frequency, Willingness) with a trait-linked action phrase. This design minimized semantic drift, ensuring alignment between output and intended trait level. In total, the SAC-Induced framework prompted each model with 11,520 questions (16 traits induced × 3 intensity levels × 5 intensity factors × 3 behavioural prompts x 16 traits observed), providing high-resolution insight into how trait intensity modulates model behavior across multiple dimensions. The prompt template in Figure 4 illustrates this setup, with annotated placeholders that have the following features:
@@
  • intensity_question: Portion of the question related to the intensity factor.
  • question: The behavioral action of the trait.
@@ 5.2.2 Results on SAC-Induced
$\Delta=\mu_{\text{induced}}-\mu_{\text{neutral}}$ (4)
@@
Figure 7: Target Trait Influence and Top Co-Movers in Mistral.
@@
Figure 5 illustrates the effect of inducing traits at intensity levels 1, 3, and 5, through deltas, allowing us to focus solely on the magnitude and direction of change. These bar charts provide a comprehensive comparison across all 16 traits for all four LLMs. As intensity increases, we observe a smooth and directional shift in trait deltas - from negative values at intensity 1, to moderate values closer to zero at intensity 3, and positive deltas at intensity 5. This trend indicates that SAC enables monotonic modulation of personality traits, with intensity 3 representing a plausible midpoint. This is a stark improvement from P2 Prompting wherein numerous negative deltas were observed even when positively keyed as seen in Tables 3 and 4. The consistency of these patterns across traits and models reinforces the robustness and semantic alignment of our framework.
@@
Next, we analyse inter-trait movement across LLMs to understand whether or not LLMs are capable of emulating latent psychological structures known in human personality science. Co-movers were selected by identifying the two traits with the largest deltas (positive or negative) accompanying the induction of a target trait as can be seen in Figure 6. While Figures 6 shows the results for Claude 3.7 Sonnet, we observed similar modulation and co-movement patterns in the other three LLMs as well, indicating that the SAC framework generalizes well across architectures (publicly accessible via Zenodo [11]).
The presence and behavior of these co-movers offer insights into whether LLMs capture inter-trait dependencies beyond the directly manipulated trait.
Table 6 summarizes the top co-movers for each target trait based on the most reactive trait shifts observed across all four LLMs. Across all experiments, we observe that all four LLMs not only respond to individual traits with fidelity but also adjust semantically related traits in a manner consistent with human personality psychology. These structured co-movements observed in Table 6 are not merely reactive but suggest that LLMs have internalized coherent personality architectures. Our intensity-control method reveals these underlying structures, and also enables precise and interpretable persona shaping.
@@
6 Limitations
While our work advances personality modelling by introducing fine-grained intensity control through SAC prompting, several limitations remain. First, the reliance on self-reported questionnaire-style responses may not fully capture deeper behavioral tendencies in LLMs. Second, observed differences in trait expression across models may conflate genuine induction effects with inherent training data distributions or architectural biases. Third, while we focus on the 16PF framework, the generalizability of our method to alternative or hierarchical personality models remains an open question for future work. Addressing these challenges will help build more robust, adaptive, and psychologically grounded AI systems.
7 Discussion and Summary of Observations
This work introduces a shift from the widely used Big Five framework to Cattell’s more granular 16 Personality Factors (16PF) model, through the construction of a new benchmark: PERS-16. Compared to the Big Five, 16PF allows for more expressive and fine-grained personality modeling, enabling the evaluation and control of traits that are typically collapsed into broader categories - such as Sensitivity and Orderliness - in traditional models. In our baseline evaluation using PERS-16, we observed that most traits naturally clustered around mid-scale values (around 3-3.5), suggesting neutral, balanced personalities. The differences across models were quantified using Euclidean distance, offering a psychometric lens for comparing LLM behavior. Applying the P2 prompting method to PERS-16 revealed its limitations. While P2 could induce some traits, its binary nature and lack of intensity control resulted in inconsistent personality shifts, often producing flat or negative deltas even for positively keyed traits. This confirmed that P2 was insufficient for controlled personality modulation within a fine-grained trait space. To overcome these issues, we proposed the Specific Attribute Control (SAC) framework, which enables trait induction across five intensity levels (1-5) along five interpretable behavioral dimensions. In SAC-neutral mode, models exhibited more stable and interpretable personality profiles than under P2. SAC-induced evaluations revealed smooth, graduated shifts in trait expression across all three LLMs. Notably, the emergence of consistent co-movers - traits that shifted in semantically or psychologically aligned ways with the target trait - highlighted the presence of latent, multidimensional personality structures within LLMs.